OpenStudy (anonymous):

3x+4=56-9x 3x+4=56-9x @Mathematics

OpenStudy (anonymous):

52=12x

OpenStudy (anonymous):

3x+9x=56-4 12x=52 x=52/12=13/3

OpenStudy (anonymous):

do the rest :)

OpenStudy (anonymous):

u do the rest

OpenStudy (anonymous):

i hve done

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

x=13/3

OpenStudy (anonymous):

atleast give the medal for it:(

OpenStudy (anonymous):

thank u naf

OpenStudy (anonymous):

no worries mighty:)